OpenStudy (tocuteforyou2):

9e-7=7e-11 solve the e @rachel.frank40 help me plz

OpenStudy (countryboyswag):

Do you know how to start this?

OpenStudy (countryboyswag):

@tocuteforyou2

OpenStudy (mathmale):

To solve 9e-7=7e-11 for the variable, " e ," combine like terms. 9e and 7e are like terms. -7 and -11 are like terms. Show your work, please.

OpenStudy (retireed):

I initial read this as exponents and said to myself, "How can I solve this, those aren't equal."

OpenStudy (tocuteforyou2):

i need help

jagr2713 (jagr2713):

Did you understand from @mathmale explanation ?

OpenStudy (mathmale):

To solve 9e-7=7e-11, begin by combining all terms that have " e " in them. Try subtracting 7e from both sides of this equation. Show your work.

OpenStudy (tocuteforyou2):

can someone give me the answer pls lol

jagr2713 (jagr2713):

It's our job as helpers to follow this website's Code of Conduct and do our best to Lead and Interact with you to get you the correct answer. Just answers won't help you learn :D

OpenStudy (mathmale):

You are given 9e-7=7e-11 Subtract 7e from both sides now, please. 9e-7=7e-11 -7e =-7e --------------- ?

OpenStudy (tocuteforyou2):

can someone help me plz

OpenStudy (tocuteforyou2):

@xXDABOIXx can you plz help me plz

OpenStudy (tocuteforyou2):

@HolsterEmission can you plz help me

OpenStudy (tocuteforyou2):

@raffle_snaffle can you help

OpenStudy (tocuteforyou2):

me plz